Find the directional derivative of f(x,y,z)=xy+z^3 at the point (2,3,1) in the direction of a vector making an angle of 3π/4 with ∇f(2,3,1).
Answer #1

Firstly , we find partial derivative

f_x=y

f_y=x

f_z=3z^2

now, we can plug points (2,3,1)

(f_x,f_y,f_z)=(3,2.3)

we are given angle between gradient and unit vector as

\theta=\frac{3\pi}{4}

so, we can find direction derivative

Direction_D=(grad-f).(u)

Direction_D=|grad-f||u|cos(\theta)

Direction_D=\sqrt{3^2+2^2+3^2}*1cos(\frac{3\pi}{4})

Direction_D=\sqrt{22}cos(\frac{3\pi}{4})

Direction_D=\sqrt{22}*\frac{-\sqrt{2}}{2}

Direction_D=\sqrt{2}*\sqrt{11}*\frac{-\sqrt{2}}{2}

Direction_D=-2*\sqrt{11}*\frac{1}{2}

Direction_D=-\sqrt{11}................ANswer